# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

### v2

@@ 7.1.1 T2I Generation and Text Editing
We also evaluate their text editing capabilities of slide images, with results illustrated in the bottom part of Fig. 19. It is observed that all models showcase unsatisfactory performance. GPT-4o partially executes the requested edits but alters portions of the original layout and modifies accompanying images. Flux.1-Kontext-dev similarly fails to locate the source text and preserve the layout. Qwen-VLo-Preview and OmniGen2 introduce extensive changes to background colors and text, producing unreadable text. BAGEL fails to respond to the instructions and leaves the image unchanged, while Janus-4o even produces chaotic content with no sense. Given rich and dense text in the slide, models struggle to locate a small portion of text while also preserving other elements.
7.2 Poster Image
7.2.1 T2I Generation and Text Editing
@@
Similar to slide images, poster stands as an important type of visually rich image that deserves investigation. The results of poster generation are shown in the top frame of Fig. 20. GPT-4o delivers the most consistent results, correctly and clearly rendering all English text, despite failing to render some small Chinese text as shown in the third row. Other models demonstrate certain abilities in text generation, while the results are often erroneous or incomplete, especially for Chinese text. Flux.1-Kontext-dev even produces a gate image instead of a poster. Models like OmniGen2, BAGEL, and Janus-4o frequently produce unreadable or nonsensical text, rendering their outputs unsuitable for practical use.
@@
In text editing, as illustrated at the bottom of Fig. 20, GPT-4o manages to correctly modify partial requested content but makes unintended changes to other parts of the poster, such as altering design elements or unrelated sections of text. Qwen-VLo-Preview partially handles Chinese text editing but often generates incomplete or incorrect modifications. The remaining models either fail to edit Chinese text altogether or produce outputs with unreadable results. None of the models, apart from GPT-4o, successfully locate and edit the correct target text in English, and all exhibit issues with altering other design elements during the editing process.
7.3 Layout-aware Text Generation
@@
Layout-aware text generation, also known as content-aware layout generation [126, 127, 128, 129, 130], requires models to position texts naturally within images, ensuring they do not oblige the original graphical component for visually appealing. The evaluation results are shown in Fig. 21. GPT-4o is the best performer that correctly adds text without compromising the original layouts in the last three rows, while it inadvertently alters the image in the first row. Qwen-VLo-Preview successfully perceives the image layout for appropriate text placement but delivers incomplete or unreadable text rendering. Other models primarily produce erroneous or readable text, and sometimes leave the images unchanged. Specifically, Flux.1-Kontext-dev suffers from illumination of generating two lines of “Camera is good”, and OmniGen2 even generates English text given a Chinese prompt (the last row). The results prove the challenge of this task for current generative models.
8 Conclusion
@@
In this paper, we comprehensively evaluate the image generation/editing capabilities of several SOTA generative models in a spectrum of OCR generative tasks. 33 common OCR tasks are chosen for evaluation, which are then categorized into five main categories based on text characteristics. We subsequently select six cutting-edge generative models, including two closed-source models, i.e., GPT-4o [57] and Qwen-VLo-Preview [60], and four open-source models, i.e., Flux.1-Kontext-dev [30] and OmniGen2 [31] (specialized generation models), as well as BAGEL [50] and Janus-4o [44] (unified understanding and generation models). Our evaluation reveals that text image generation and editing remain significant challenges for current up-to-date models, in which they suffer from inaccurate text locating, poor structural preservation, blurred or illegible character generation, etc. Given the lack of dedicated evaluation for image generation performance in the OCR field, our work fills this gap for the first time and provides crucial insights into current generative models’ limitations and potential future directions. We hope our analysis can provide useful insights for the community and inspire increased efforts to improve text image synthesis and editing, internalizing high-quality text image generation skills into general-domain generative models to advance our step toward AGI.

### v3

@@ 7.1.2. Text Editing
We also evaluate their text editing capabilities of slide images, with results illustrated in the bottom part of Fig. 19. The overall performance is unsatisfactory. GPT-4o partially applies edits but disrupts the original layout and modifies accompanying images. Flux.1-Kontext-dev similarly fails to locate the source text and preserve the layout. Qwen-Image modifies the wrong text. Qwen-VLo-Preview and OmniGen2 alter backgrounds and text extensively, yielding unreadable results. BAGEL fails to respond to the instructions and leaves the image unchanged, while Janus-4o even produces chaotic and nonsensical content. Given rich and dense text in the slide, models struggle to locate a small portion of text while also preserving other elements.
7.2. Poster Image
7.2.1. T2I Generation
@@
Similar to slide images, poster stands as an important type of visually rich image worth studying. Results of poster generation are shown in the top frame of Fig. 20. GPT-4o delivers the most consistent results, correctly and clearly rendering all English text, despite failing to render some small Chinese text (the third row). Qwen-Image is the only one that correctly generates small Chinese text, while the second generated poster somewhat lacks visual appeal. Other models demonstrate certain abilities in text generation, with frequent errors or omissions, especially in Chinese. Flux.1-Kontext-dev even produces a gate image instead of a poster. Models like OmniGen2, BAGEL, and Janus-4o frequently produce unreadable or nonsensical text, rendering their outputs unsuitable for practical use.
7.2.2. Text Editing
@@
Results of text editing illustrated at the bottom of Fig. 20, GPT-4o correctly modifies partial requested content but also unexpectedly alters unrelated design elements. Qwen-VLo-Preview partially handles Chinese text editing but generates incomplete or incorrect modifications, while Qwen-Image misplaces text and produces unreadable content. Remaining models either fail to edit Chinese text or produce outputs with unreadable results. Except for GPT-4o, none successfully locate and edit the correct English text, and all introduce unintended changes to other elements. This indicates significant room for improvement on this task, particularly for open-source models.
7.3. Layout-aware Text Generation
@@
Layout-aware text generation, also known as content-aware layout generation (Hsu et al., 2023; Horita et al., 2024; Seol et al., 2025; Guo et al., 2025), requires placing text naturally within images without obliging the original graphical component. Results are shown in Fig. 21. GPT-4o performs best, which correctly adds text without compromising the original layouts in most cases, although it may inadvertently alter the image (the first row). Qwen-VLo-Preview preserves layout but often renders text incompletely. Other models primarily produce erroneous or unreadable text, or sometimes leave the images unchanged. Specifically, Flux.1-Kontext-dev redundantly generates two lines of “Camera is good”, Qwen-Image hallucinates human hands (the last row), and OmniGen2 generates English text given a request for Chinese (the last row). These results highlight the challenge of this task for these models.
8. Conclusion
@@
In this paper, we comprehensively evaluate the image generation/editing capabilities of several SOTA generative models across a spectrum of OCR generative tasks. We select 33 common OCR tasks and categorize them into five main groups based on text characteristics. Seven cutting-edge generative models are then evaluated, including two closed-source models, i.e., GPT-4o (OpenAI, 2025) and Qwen-VLo-Preview (Team, 2025) (unified generative models), and five open-source models, i.e., Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), and OmniGen2 (Wu et al., 2025d) (specialized generation models), as well as BAGEL (Deng et al., 2025) and Janus-4o (Chen et al., 2025b) (unified generative models). Our evaluation reveals that text image generation and editing remain significant challenges for current models, in which they suffer from inaccurate text locating, poor structural preservation, blurred or illegible character generation, etc. By providing the first dedicated evaluation of OCR-focused image generation, our work not only exposes the limitations of current generative models but also points out potential directions for future development. We envision that high-quality text image generation and editing could be internalized into general-domain generative models as foundational skills, and we hope our analysis inspires the community to advance toward this goal.